OpenStudy (anonymous):

is average rate of change of the function, g(x)=x^2, over interval [-1,1] equal to 0?

OpenStudy (amistre64):

the average rate of change over any interval [a,b] is just: f(b) - f(a) -------- b-a

OpenStudy (amistre64):

in your case: \[\frac{g(1)-g(-1)}{1-(-1)}\]

myininaya (myininaya):

\[\frac{(1)^2-(-1)^2}{1-(-1)}=\frac{1-1}{1+1}=\frac{0}{2}=0\]

OpenStudy (amistre64):

and yes, the average rate of change over a parabola between 2 points that are the same height is 0 :)

OpenStudy (anonymous):

Thank you!